Question 1 Skysong Company purchased equipment for $270,000 on October 1, 2017. It is estimated that the equipment will have a useful life of 8 years and a salvage value of $15,000. Estimated production is 40,800 units and estimated working hours are 19,700. During 2017, Skysong uses the equipment for 520 hours and the equipment produces 1,000 units Compute depreciation expense under each of the following methods. Skysong is on a calendar-year basis ending December 31. (Round rate per hour and rate per unit to 2 decimal places, e.g. 5.35 and final answers to 0 decimal places, e.g. 45,892.) (a) Straight-line method for 2017 (b) Activity method (units of output) for 2017 (c) Activity method (working hours) for 2017 (d) Sum-of-the-years-digits method for 2019 (e) Double-declining-balance method for 2018 Click if you would like to Show Work for this question: Open Show Work
Answer #1

Ans:

S.NO Method Depreciation
a Straight line method for 2017 $7,969
b Activity method (Units of Output) for 2017 $6,250
c Activity method (Working Hours) for 2017 $6,731
d

Sum-of the-years Digit method for 2019:

$47,812
e Double-declining balance method for 2018 $63,281

Working:

a)Straight line method for 2017 :

cost of the equipment – scrap value / useful life

Cost of Equipment =$270,000

Salvage Value = $15,000

Useful economic life = 8 years

Straight line depreciation per year: ($270,000 - $15,000) / 8 = $ 31,875

Depreciation for the year ended December 2017 = $ 31,875 / 12 * 3 = $7,969 (Approx.)

b)Activity method (Units of Output) for 2017 :

Estimated production units = 40,800 for the whole life

During 2017 equipment produces 1000 units

Depreciation for the year 2017: ($270,000 - $15,000) / 40,800 units * 1000 units = $6,250

c)Activity method (Working Hours) for 2017:

Estimated production hours = 19,700 hours

During 2017 production worked for 520 hours

Depreciation for the year 2017: ($270,000 - $15,000) / 19,700 hours * 520 hours = $6,731

d) Sum-of the-years Digit method for 2019:

First year of 8 year life = 8 / (8+7+6+5+4+3+2+1) = 8/36

For Jan 19 to Sept 19 = 7/36 for 9 months

For Oct 19 to Dec 19 = 6/36 for 3 months

Depreciation for the year 2019: {($270,000 - $15,000) * 7/36 * 9/12} + {($270,000 - $15,000) * 6/36 * 3/12}

= $37,187+10,625 = $47,812

e)Double-declining balance method for 2018 :

Straight line depreciation would be 1/8 or 12.5% p.a.

Double declining depreciation: 12.5% * 2 = 25% p.a

Written Down Value as on Jan 1, 2018 = $270,000 – ($270,000 * 25% * 3/12) = $253,125

Depreciation for the year 2018: 25% of $253,125 = $63,281